# Mutual Information Free Topological Generalization Bounds via Stability

### v1

@@ C.1 Implementation Details
We implement training and validation loops in PyTorch Lightning Version $2.4$ using pytorch version $2.4.1$. Both models are trained using the AdamW optimizer with a cosine annealing learning rate scheduler.
@@
The ViT is configured for image classification on the CIFAR100 dataset. For the Vision Transformer on CIFAR100, we utilize a model with approximately 1.2 million parameters. This represents a reduction in size compared to the ViT experiments in [ADS+24], as the empirical analyses we conduct require substantially more experiments due to the expansion of training set size $n$ and thus storage of weight iterates.
@@
In addition to the main paper results, we analyze GraphSAGE [HYL17] on the graph classification task for MNISTSuperpixels [MBM+17]. It processes graphs where each node has a single input feature. These features are initially projected to a hidden dimension of 128. For graph-level predictions, node embeddings are aggregated using ‘mean’ global pooling, followed by a linear layer for classification into 10 classes. The model uses 2-dimensional positional encodings (representing x,y coordinates), which are added to the node features. We use 4 GraphSAGE convolutional layers with a dropout rate of 0.05, a base learning rate of 0.001, and weight decay of $0.0005$ to achieve an adequate generalization gap.
@@ C.2 Stability Analysis
We first elaborate on our procedure to estimate trajectory stability in Section 4.1. The overall algorithm, a proxy to asses 3.1, is summarized in Algorithm 1. Stability experiments were performed with a fixed learning rate ($\eta$) and batch size ($b$), while varying random seeds, as described in Section 4.1. To ensure we capture two distinct training regimes, we analyze both randomly initialized and locally converged models; the latter is standard for the analysis of topological invariants [ADS+24]. We train each model for 500 additional iterations when analyzing trajectory stability.
The same setup is then conducted twice: once with the replacement set $J$ Section 4.1, and once without. Trajectory stability (3.1) is then estimated through the Hausdorf distance over each pair of trajectories where all model parameters are fixed except for the replacement set $J$.
@@
1:Training set $S=\{z_{1},\dots,z_{n}\}$
2:Injection set $S_{J}=\{z^{\prime}_{1},\dots,z^{\prime}_{J}\}$ with $S_{J}\cap S=\emptyset$, and $J\leq n$
3:Estimate of stability parameter:
$\max_{w\in\mathcal{W}_{S,U}}\min_{w^{\prime}\in\mathcal{W}_{S^{\prime},U}}\sup_{Z\in\mathbf{Z}}|\ell(w,Z)-\ell(w^{\prime},Z)|$
4:procedure Training Phase
5:     Sample random indices $I=\{i_{1},\dots,i_{J}\}\subset\{1,\dots,n\}$
6:     Replace $\{z_{i}:i\in I\}$ with $S_{J}$ to form new dataset:
$S_{\text{new}}:=\left(S\setminus\{z_{i}:i\in I\}\right)\cup S_{J}$
7:     Train $T$ iterations on $S$ to obtain $\mathcal{W}_{S,U}=\{w_{1},\dots,w_{T}\}$
8:     Train $T$ iterations on $S_{\text{new}}$ to obtain $\mathcal{W}_{S_{\text{new}},U}=\{w^{\prime}_{1},\dots,w^{\prime}_{T}\}$
9:procedure Estimation Phase
10:     Set $M:=\min\{n,500\}$
11:     Sample $M$ indices $L=\{j_{1},\dots,j_{M}\}\subset\{1,\dots,n\}$
12:     Form evaluation set:
$\mathbf{Z}:=\{z_{j}\in S_{\text{new}}:j\in L\}\setminus S_{J}$
13:     Let $N:=|\mathbf{Z}|=M-J$
14:     Initialize matrices $A^{(1)},\dots,A^{(N)}\in\mathbb{R}^{T\times T}$
15:     for $i=1$ to $N$ do
16:         for $j=1$ to $T$ do
17:              for $l=1$ to $T$ do
18:                  $A^{(i)}_{j,l}:=|\ell(w_{j},Z_{i})-\ell(w^{\prime}_{l},Z_{i})|$
19:     Define $A\in\mathbb{R}^{T\times T}$ by:
$A_{j,l}:=\max_{i\in\{1,\dots,N\}}A^{(i)}_{j,l}$
20:     Compute vector $a\in\mathbb{R}^{T}$ where:
$a_{j}:=\min_{1\leq l\leq T}A_{j,l}$
21:     return Final estimate: $\max_{1\leq j\leq T}a_{j}$
C.3 Topological Complexity Analysis
@@
Following [ADS+24], we fine-tune each model from a local minimum for an additional 5000 iterations after training to convergence. We use a reduced parameter grid of $4\times 4$ for $\eta$ and $b$ ([ADS+24] use $6\times 6$), due to computational constraints arising from the varying the training set size $n$. Detailed hyperparameters are listed in Table 1.
@@
C.4 Hardware and Runtime
Models are trained on a cluster with 40GB A40/A100 GPUs; although the experiments are reproducible with lower memory requirements. For our grid experiments, computing topological features alongside hyperparameter variations (sample size, learning rate, and batch size) demanded approximately 57 GPU hours for GraphSage and 140 GPU hours for the Vision Transformer (ViT). In contrast, the stability analysis on the ViT required about 50 total GPU hours.
@@ Appendix D Additional Empirical Results
In this section, we present additional empirical analyses and further substantiate the main paper analysis in Section 4.1 and Section 4.1.
D.1 Stability Estimation
In this section, we continue with the analysis of Section 4.1. We observe that, for both seen and unseen samples, an increase in sample size leads to a decrease in the estimated stability parameter (see Table 2). As shown in Section 4.1 and Table 2, the locally converged model demonstrates comparable stability between seen and unseen samples, with values similar to those of the randomly initialized model. This empirical evidence suggests that our trajectory stability assumptions are both theoretically sound and practically achievable in real-world applications.
Pretrain States Loss Difference Dataset Size ($n$) 100 500 1000 5000 10000 Randomly Initialized Seen 1.23 ± 0.22 0.56 ± 0.07 0.32 ± 0.04 0.07 ± 0.01 0.04 ± 0.01 Unseen 1.39 ± 0.16 0.60 ± 0.08 0.32 ± 0.04 0.07 ± 0.01 0.04 ± 0.00 Locally Converged Seen 1.20 ± 0.14 2.76 ± 0.38 2.37 ± 0.32 0.37 ± 0.07 0.36 ± 0.10 Unseen 4.60 ± 0.13 2.97 ± 0.23 1.59 ± 0.10 0.51 ± 0.10 0.28 ± 0.04
Table 2: Stability Analysis of Loss Differences Across Pretraining Iterations and Dataset Sizes for ViT Cifar100 (1.2M).
D.2 Topological Complexity Analysis
Building on the empirical analysis in Section 4.1, we elaborate further on how our topological complexities evolve with dataset size $n$.
Vision Transformer
Table 3 reports topological quantities for multiple scaling parameters of the positive magnitude, as well as $\mathbf{E}^{1}$. We observe a consistent trend across all topological features, mirroring the patterns identified in Section 4.1. This suggests that the observed behavior is not isolated to a specific topological quantity.
GraphSage
In addition to the experiments on the Vision Transformer, we validate our observation on a conceptually different model. We conduct the same experiment for GraphSage, as introduced in Section C.1. In these experiments, we similarly observe an increase in the magnitude of topological complexities with increasing sample size $n$. Figure 5 depicts a representative example.
While a similar behavior of increasing slope is apparent for GraphSage, the overall pattern is less apparent. This can be partially attributed to the larger discrepancies in the generalization gap, which are more difficult to modulate precisely by varying $\eta$ and $b$.
Nevertheless, as seen in the relationship between the generalization gap and $\mathbf{E}^{1}$ Figure 5, the slope similarly trends upwards with sample size $n$.
We additionally report positive magnitude for scaling that is theoretically supported ($3]{n}$), and the fixed scale $\mathbf{PMag}(100\cdot\mathcal{W_{S}})$. For the theoretically supported scales of $\mathbf{PMag}(\sqrt[3]{n}\cdot\mathcal{W_{S}})$, we observe a similar tendency of an increase in slope. Because the positive magnitude intuitively represents the effective number of distinct points in a space at a given scale $\lambda$, excessively large values of $\lambda$ can produce distorted and uninformative quantities. This occurs as fine-grained structural details are lost when the scale exceeds meaningful topological variations.
Figure 4: Topological invariants $\mathbf{E}^{1}$ and PMag for GraphSage on MNISTSuperpixel. PMag is depicted for the fixed scale of 100. The mean and std. err. over the batch sizes are shown for learning rates $10^{-4},10^{-3}$.
Figure 5: Behavior of $\mathbf{E}^{1}$ over $n$ for GraphSage on MNISTSuperpixel. Pearson correlation is reported per training subset $n$.
ViT Cifar100 Measure Dataset Size ($n$) 100 500 1000 5000 10000 $\mathbf{PMag}(\sqrt[3]{100}\cdot\mathcal{W_{S}})$ $\tau$ 0.733 0.767 0.667 0.717 0.633 $r$ 0.971 0.914 0.912 0.928 0.822 Slope 47.225 97.009 172.533 384.941 511.956 $\mathbf{PMag}(\sqrt[3]{500}\cdot\mathcal{W_{S}})$ $\tau$ 0.717 0.800 0.700 0.767 0.633 $r$ 0.973 0.921 0.924 0.940 0.834 Slope 87.332 207.714 371.531 790.149 1043.039 $\mathbf{PMag}(\sqrt[3]{1000}\cdot\mathcal{W_{S}})$ $\tau$ 0.717 0.817 0.700 0.750 0.633 $r$ 0.974 0.930 0.931 0.944 0.838 Slope 113.900 278.563 493.025 1021.470 1342.187 $\mathbf{PMag}(\sqrt[3]{5000}\cdot\mathcal{W_{S}})$ $\tau$ 0.700 0.800 0.733 0.767 0.650 $r$ 0.975 0.952 0.944 0.948 0.843 Slope 205.402 499.767 851.550 1655.984 2151.617 $\mathbf{PMag}(\sqrt[3]{10000}\cdot\mathcal{W_{S}})$ $\tau$ 0.700 0.783 0.733 0.750 0.633 $r$ 0.975 0.960 0.947 0.947 0.842 Slope 259.753 616.699 1030.246 1948.297 2519.328 $\mathbf{E}^{\alpha}$ $\tau$ 0.750 0.750 0.650 0.683 0.533 $r$ 0.976 0.957 0.942 0.948 0.842 Slope 33.652 73.124 123.392 247.425 325.163 $\mathbf{PMag}(100\cdot\mathcal{W_{S}})$ $\tau$ 0.783 0.767 0.650 0.700 0.650 $r$ 0.947 0.981 0.921 0.893 0.794 Slope 814.032 1400.576 2013.602 3177.603 4022.352
Table 3: We report the Kendall correlation coefficient $\tau$, the Pearson correlation coefficient $r$, and the slope for the Vision Transformer model on Cifar100. Bolded values denote the theoretically optimal scaling parameter of $\mathbf{PMag}$ corresponding to sample size n.
GraphSage Measure Dataset Size ($n$) 100 500 1000 5000 10000 $\mathbf{PMag}(\sqrt[3]{100}\cdot\mathcal{W_{S}})$ $\tau$ 0.850 0.533 0.683 0.667 0.500 $r$ 0.896 0.682 0.924 0.566 0.233 Slope 39.653 62.366 48.664 124.134 75.402 $\mathbf{PMag}(\sqrt[3]{500}\cdot\mathcal{W_{S}})$ $\tau$ 0.867 0.617 0.750 0.667 0.517 $r$ 0.897 0.691 0.929 0.564 0.241 Slope 67.973 107.236 79.759 226.185 142.980 $\mathbf{PMag}(\sqrt[3]{1000}\cdot\mathcal{W_{S}})$ $\tau$ 0.900 0.617 0.767 0.717 0.533 $r$ 0.898 0.692 0.928 0.570 0.250 Slope 85.249 134.828 96.132 276.827 180.090 $\mathbf{PMag}(\sqrt[3]{5000}\cdot\mathcal{W_{S}})$ $\tau$ 0.883 0.583 0.800 0.683 0.517 $r$ 0.904 0.691 0.914 0.598 0.280 Slope 142.404 225.500 140.712 403.594 283.260 $\mathbf{PMag}(\sqrt[3]{10000}\cdot\mathcal{W_{S}})$ $\tau$ 0.883 0.583 0.800 0.683 0.467 $r$ 0.908 0.690 0.900 0.617 0.298 Slope 176.644 279.671 163.426 461.667 335.112 $\mathbf{E}^{\alpha}$ $\tau$ 0.883 0.550 0.700 0.483 0.367 $r$ 0.921 0.683 0.922 0.571 0.279 Slope 19.361 30.711 26.662 78.674 56.366 $\mathbf{PMag}(100\cdot\mathcal{W_{S}})$ $\tau$ 0.917 0.567 0.767 0.633 0.483 $r$ 0.942 0.696 0.729 0.796 0.541 Slope 635.358 1092.277 433.767 941.365 855.726
Table 4: We report the Kendall correlation coefficient $\tau$, the Pearson correlation coefficient $r$, and the slope for the GraphSage model on MNISTSuperpixel. Bolded values denote the theoretically optimal scaling parameter of $\mathbf{PMag}$ corresponding to sample size n.

### v2

@@ C.1 Implementation Details
We implement training and validation loops in PyTorch Lightning Version $2.4$ using pytorch version $2.4.1$. Both models are trained using the AdamW optimizer with a cosine annealing learning rate scheduler. The training was stopped after reaching an accuracy of $95\%$. No regularization techniques were used to prevent confounding factors in each fine-tuning task.
@@
The ViT is configured for image classification on the CIFAR100 dataset. For the Vision Transformer on CIFAR100, we utilize a model with approximately 1.2 million parameters. This represents a reduction in size compared to the ViT experiments in Andreeva et al. (2024), as the empirical analyses we conduct require substantially more experiments due to the expansion of training set size $n$ and thus storage of weight iterates.
@@
We analyze GraphSAGE Hamilton et al. (2017) on the graph classification task for MNISTSuperpixels Monti et al. (2017). This model consists of approximately 0.2 million parameters. It processes graphs where each node has a single input feature. These features are initially projected to a hidden dimension of 128. For graph-level predictions, node embeddings are aggregated using ‘mean’ global pooling, followed by a linear layer for classification into 10 classes. The model uses 2-dimensional positional encodings (representing x,y coordinates), which are added to the node features. We use 4 GraphSAGE convolutional layers with a dropout rate of 0.05, a base learning rate of 0.001, and weight decay of $0.0005$ to achieve an adequate generalization gap.
@@ C.2 Stability Analysis
We first elaborate on our procedure to estimate random set stability in Section˜5. The overall algorithm, a proxy to asses Section˜5, is summarized in Algorithm˜1. Stability experiments were performed with a fixed learning rate ($\eta$) and batch size ($b$), while varying random seeds (e.g. $\{100,200,300,400,500\}$. The same setup is then conducted twice: once with the replacement set of size $J=50$, and once without. Since ˜3.1 requires taking the supremum over all samples $Z\in\mathcal{Z}$, we approximate it by evaluating the stability parameter on 500 samples seen during training and 500 unseen samples. In Table˜1, we report the maximum value across both sets.
@@
[…]
5:procedure Procedure(Training)
6:  Let $I=\{i_{1},...,i_{J}\}$ be an uniformly random subset of $\{1,...,n\}$ $\triangleright$ Indices to replace
7:  $S_{new}:=S\setminus\{z_{i}:i\in I\}\cup S_{J}:=\{z_{new,1},...,z_{new,n}\}$
8:  Generate $\mathcal{W}_{S,U}$ with $|\mathcal{W}_{S,U}|=T$
9:  Generate $\mathcal{W}_{S_{new},U}$ with $|\mathcal{W}_{S_{new},U}|=T$
10:procedure Procedure(Estimation)
11:  $M=\min\{n,500\}$
12:  $L=\{j_{1},...,j_{M}\}$ be an uniform random subset of $\{1,...,n\}$
13:  $S_{eval}:=\{z_{new,i}:i\in L\}\setminus S_{J}$
14:  $N:=|S_{eval}|=M-J$
15:
16:  Initialize $A^{1},...,A^{N}\in\mathbb{R}^{T\times T}$
17:  Let $w_{l}\in\mathcal{W}_{S,U}\ \forall l\in\{1,...,T\}$
18:  Let $w_{new,l}\in\mathcal{W}_{S_{new},U}\ \forall l\in\{1,...,T\}$
19:  Let $\ell(\cdot,\cdot)$ denote the used loss function.
20:
21:  For all $i\in\{1,...,N\}$ compute
22:  for $j\leftarrow 1$ to $T$ do
23:   for $l\leftarrow 1$ to $T$ do
24:     $A^{i}_{j,l}:=|\ell(w_{j},Z_{i})-\ell(w_{new,l},Z_{i})|$ $\triangleright$ $Z_{i}\in S_{eval}$
25:  Now we compute $A:=\frac{1}{N}\sum_{i=1}^{N}A^{i}$
26:
27:  Now we take the minimum over each row in matrix AA
28:  Initialize $a\in\mathbb{R}^{T}$ and compute
29:  for $m\leftarrow 1$ to $T$ do $a_{m}=\min_{1\leq l\leq T}A_{m,l}$
30:  The algorithm ends with $\max_{1\leq h\leq m}a_{h}$
C.3 Bound Estimation
In the first part of Section˜5, we estimate the Rademacher complexity bound stated in Lemma˜3.4. In particular, we optimize the quantity
$2\sqrt{\tfrac{2\log T}{J}}\;+\;2J\beta_{n},$
which is minimized at
$J^{*}\;=\;2^{-\tfrac{1}{3}}\,\beta_{n}^{-\tfrac{2}{3}}\,(\log T)^{-\tfrac{1}{3}}.$
Since $J$ must be a divisor of the sample size $n$ (see Lemma˜3.4), we define the effective choice of $J$ as
$J\;=\;\underset{d\mid n}{\arg\min}\,\big|d-J^{*}\big|.$
C.4 Topological Complexity Analysis
@@
Following Andreeva et al. (2024), we fine-tune each model from a local minimum for an additional 5000 iterations after training to convergence. We use a reduced parameter grid of $4\times 4$ for $\eta$ and $b$ (Andreeva et al. (2024) use $6\times 6$), due to computational constraints arising from the varying the training set size $n$. Detailed hyperparameters are listed in Table˜2.
@@
C.5 Hardware and Runtime
Models are trained on a cluster with 40GB A40/A100 GPUs; although the experiments are reproducible with lower memory requirements. For our grid experiments, computing topological features alongside hyperparameter variations (sample size, learning rate, and batch size) demanded approximately 57 GPU hours for GraphSage and 140 GPU hours for the Vision Transformer (ViT). In contrast, the stability analysis required about 50 total GPU hours.
@@ Appendix D Additional Empirical Results
In this section, we present additional empirical results. The first part complements the second part of Section˜5. In the main text, we focused on the $\alpha$-weighted lifetime sum in Theorem˜4.4. The second bound, given in terms of the positive magnitude $\mathbf{PMag}$ will be discussed here.
Motivated by the dependence of the stability parameter in ˜3.1 on the sample size $n$, we also investigate the topological complexity measures introduced in Theorem˜4.4. To the best of our knowledge, this behavior has not been addressed in previous work. We study it here to provide a more complete picture of the $\alpha$-weighted lifetime sum and the positive magnitude.
D.1 Interplay between Stability and $\mathbf{C}(\mathcal{W}_{S,U})$
In this subsection, in addition to the results presented in Section˜5, we provide the results specifically for the positive magnitude. The second bound in Theorem˜4.4 depends on the scaling parameter $s(\lambda)$. Recall that in Theorem˜4.4, the parameter $s(\lambda)$ can be chosen arbitrarily. The theoretically justified choice is $s(\lambda)=\mathcal{O}(n^{-1/3})$ in order to obtain the optimal convergence rate. Nevertheless, alternative choices of $s(\lambda)$ have also led to remarkable correlation results in previous work Andreeva et al. (2024).We focus on the setting $s(\lambda)=n^{-1/3}$ with $n\in\{100,10^{4}\}$.
Results and Discussion
We observe the same behavior as in the main text (Section˜5). We note that the effects for GraphSage Figure˜5 and Figure˜5 are less pronounced than for the ViT (Figure˜7 and Figure˜7); nonetheless, the general trend remains visible. Interestingly, for $s(\lambda)=0.01$, we observe a different behavior. While for higher values of $s(\lambda)$ the steepness varies with the sample size, as predicted by Theorem˜4.4 and also observed for $\mathbf{E}^{1}$, the steepness remains consistent across different values of $n$.
The same behavior for $\mathbf{PMag}(s(\lambda)\cdot\mathcal{W}_{S,U})$, with $s(\lambda)=n^{\frac{1}{3}}$, can be explained theoretically. In Theorem˜4.4, we observe the product structure. Indeed, Theorem˜4.4 asserts that $\log\mathbf{E}^{1}(\mathcal{W}_{S,U})$ should be approximately of order at least $\beta_{n}^{-1/3}G_{S}(\mathcal{W}_{S,U})$, which corresponds to $n^{1/3}G_{S}(\mathcal{W}_{S,U})$ when $\beta_{n}=\Theta(1/n)$. Therefore, our experimental results provide strong empirical support for Theorem˜4.4 both for $\mathbf{E}^{1}$ and $\mathbf{PMag}$.
Figure 4: Variation of $\mathbf{PMag}(\sqrt[3]{100}\cdot\mathcal{W}_{S,U})$ with sample size $n$ for the model GraphSage. Pearson correlation coefficients $r$ are reported for each subgroup.
Figure 5: Variation of $\mathbf{PMag}(\sqrt[3]{10000}\cdot\mathcal{W}_{S,U})$ with sample size $n$ for the model GraphSage. Pearson correlation coefficients $r$ are reported for each subgroup.
Figure 6: Variation of $\mathbf{PMag}(\sqrt[3]{100})$ with sample size $n$ for the model ViT. Pearson correlation coefficients $r$ are reported for each subgroup.
Figure 7: Variation of $\mathbf{PMag}(\sqrt[3]{10000})$ with sample size $n$ for the model ViT. Pearson correlation coefficients $r$ are reported for each subgroup.
D.2 Complexity Measures $\mathbf{C}(\mathcal{W}_{S,U})$ as functions of the sample size $n$
While the main write-up in Section˜5 focused on order of the bound values of Lemma˜3.4 and overall behavior, in this section we investigate the behavior of the topological complexity measure introduced in Andreeva et al. (2024). In particular, in Theorem˜4.4 we recover the same complexity measures $\mathbf{C}(\mathcal{W}_{S,U})$ within our stability framework. Motivated by the dependence of the stability parameter in ˜3.1 on the sample size $n$, we investigate the positive magnitude and the $\alpha$-weigthed lifetime sum with respeect to the training data size.
Experimental Design
Similarly to Section˜5, we investigate the behavior of two conceptually different models: a Vision Transformer (ViT) and GraphSage. The implementation details remain consistent with the main experimental setting. In particular, we analyze the behavior of $\mathbf{E}^{1}(\mathcal{W}_{S,U})$ as a function of the sample size $n$, for fixed learning rates $\eta=10^{-4}$ and $\eta=10^{-5}$ for ViT, and $\eta=10^{-3}$ and $\eta=10^{-4}$ for GraphSage, across all batches. Our focus is on $\mathbf{E}^{1}(\mathcal{W}_{S,U})$ and $\mathbf{PMag}(s(\lambda)\cdot\mathcal{W}_{S,U})$, with $s(\lambda)=n^{\frac{1}{3}}$ and $n\in\{100,10^{4}\}$.
Results and Discussion
For fixed learning rates, we observe that both $\mathbf{E}^{1}$ and $\mathbf{PMag}$ increase with increasing $n$ and appear to stabilize for large values of $n$ (see Figures 9, 9, 11, 11, 13, 13 . This suggests a relationship between the sample size and the topology of the training trajectory, a phenomenon that has not been previously highlighted in the literature.
At first glance, the increase of $\mathbf{E}^{1}(W_{S},U)$ with $n$ may seem counterintuitive, since we typically expect generalization error (and related bounds) to decrease as the sample size grows. However, this behavior is consistent with theory: the $\alpha$-weigthed lifetime sum is expected to grow at most polynomially with $n$. The positive magnitude remains bounded in terms of the number of iterations.
Figure 8: Evolution of $\mathbf{E}^{1}(\mathcal{W}_{S,U})$ with increasing sample size n, for learning rates $\eta=10^{-5}$ and $\eta=10^{-4}$ (ViT).
Figure 9: Evolution of $\mathbf{E}^{1}(\mathcal{W}_{S,U})$ with increasing sample size n, for learning rates $\eta=10^{-3}$ and $\eta=10^{-4}$ (GraphSage).
Figure 10: Evolution of $\mathbf{PMag}(\sqrt[3]{100}\cdot\mathcal{W}_{S,U})$ with increasing sample size n, for learning rates $\eta=10^{-5}$ and $\eta=10^{-4}$ (ViT).
Figure 11: Evolution of $\mathbf{PMag}(\sqrt[3]{100}\cdot\mathcal{W}_{S,U})$ with increasing sample size n, for learning rates $\eta=10^{-3}$ and $\eta=10^{-4}$ (GraphSage).
Figure 12: Evolution of $\mathbf{PMag}(\sqrt[3]{10000}\cdot\mathcal{W}_{S,U})$ with increasing sample size n, for learning rates $\eta=10^{-4}$ and $\eta=10^{-4}$ (ViT).
Figure 13: Evolution of $\mathbf{PMag}(\sqrt[3]{10000}\cdot\mathcal{W}_{S,U})$ with increasing sample size n, for learning rates $\eta=10^{-3}$ and $\eta=10^{-4}$ (GraphSage).